# OWSM-CTC: An Open Encoder-Only Speech Foundation Model for Speech Recognition, Translation, and Language Identification

### v1

@@ Abstract
There has been an increasing interest in large speech models that can perform multiple speech processing tasks in a single model. Such models usually adopt the encoder-decoder or decoder-only architecture due to their popularity and good performance in many domains. However, autoregressive models can be slower during inference compared to non-autoregressive models and also have potential risks of hallucination. Though prior studies observed promising results of non-autoregressive models for certain tasks at small scales, it remains unclear if they can be scaled to speech-to-text generation in diverse languages and tasks. Inspired by the Open Whisper-style Speech Model (OWSM) project, we propose OWSM-CTC, a novel encoder-only speech foundation model based on Connectionist Temporal Classification (CTC). It is trained on 180k hours of public audio data for multilingual automatic speech recognition (ASR), speech translation (ST), and language identification (LID). Compared to encoder-decoder OWSM, our OWSM-CTC achieves competitive results on ASR and up to 25% relative improvement on ST, while it is more robust and 3 to 4 times faster for inference. OWSM-CTC also improves the long-form ASR result with 20x speed-up. We will publicly release our codebase, pre-trained model, and training logs to promote open science in speech foundation models.
@@ 1 Introduction
111122223333444488881010101012121212mediumbaseSpeed-up ($\rightarrow$)WER ($\leftarrow$)
@@
1111222233334444101010101515151520202020252525253030303035353535mediumbaseSpeed-up ($\rightarrow$)BLEU ($\rightarrow$)
@@
1111222233334444555510101010151515152020202025252525mediumbaseSpeed-up ($\rightarrow$)BLEU ($\rightarrow$)
@@
The great success of large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023; Anil et al., 2023b) has sparked a growing interest in developing foundation models in various modalities. Recent studies have explored different approaches towards multilingual and multi-tasking speech foundation models (Radford et al., 2023; Zhang et al., 2023; Pratap et al., 2023; Rubenstein et al., 2023; Barrault et al., 2023; Peng et al., 2023e). OpenAI’s Whisper (Radford et al., 2023) is a series of Transformer encoder-decoder models trained on 680k hours of proprietary labelled audio. Whisper achieves strong results in multilingual automatic speech recognition (ASR), any-to-English speech translation (ST), and spoken language identification (LID). Although it shows the effectiveness of large-scale (weakly) supervised pre-training, the full development pipeline including training data details is not publicly accessible. Recent work releases Open Whisper-style Speech Models (OWSM) (Peng et al., 2023e, 2024) with the aim of reproducing Whisper-style training using public data and open-source toolkits. However, Whisper and OWSM adopt the encoder-decoder architecture, which generates text tokens given speech in an autoregressive manner. They might hallucinate during inference, and the speed can be slow. Other models with a decoder-only architecture like AudioPaLM (Rubenstein et al., 2023) and VioLA (Wang et al., 2023b) would suffer from the same issues due to autoregressive decoding.
@@
Another type of work like Google’s USM (Zhang et al., 2023) and Meta’s MMS (Pratap et al., 2023) uses non-autoregressive models with Connectionist Temporal Classification (CTC) Graves et al. (2006), but these CTC-based models are designed for ASR only. Prior studies have also achieved promising results of CTC models for ST only, but they mainly focus on specific language pairs at much smaller scales (Inaguma et al., 2021; Chuang et al., 2021; Xu et al., 2023). Some of them employ additional decoders (Inaguma et al., 2021; Yan et al., 2023) or cross-attention layers (Xu et al., 2023), making the model more complicated.
@@
A natural question now arises: Can we build a non-autoregressive encoder-only model for speech-to-text generation in diverse languages and multiple tasks like Whisper/OWSM? This research problem has become increasingly important in the era of LLMs, because large-scale pre-trained speech encoders can serve as an adaptor between the speech and text modalities (Gong et al., 2023; Wang et al., 2023a), providing a promising avenue towards general-purpose multi-modal foundation models (Anil et al., 2023a).
@@
In this work, we propose OWSM-CTC, a novel encoder-only speech foundation model based on multi-task self-conditioned CTC (Nozaki and Komatsu, 2021) to imitate OWSM’s multilingual ASR, any-to-any ST, and LID functionalities. Following previous encoder-decoder OWSM v3.1 models (Peng et al., 2024), we train a 1B OWSM-CTC model using 180k hours of public data covering 151 languages. Extensive evaluations show that our OWSM-CTC exhibits strong performance and efficiency. Compared to the 1B OWSM v3.1 medium model, OWSM-CTC achieves comparable performance for ASR and superior performance for various ST directions (up to 25% relative improvement) while being more robust and showing 3 to 4 times inference speed-up. For long-form ASR, OWSM-CTC improves the WER and is 20 times faster due to the batched parallel decoding. OWSM-CTC also outperforms the other models on LID. We will publicly release our codebase, pre-trained model weights, and training logs to facilitate the development of large speech models.
@@ 2.1 Speech foundation models
Attention-based encoder-decoder. OpenAI’s Whisper (Radford et al., 2023) adopts the standard Transformer encoder-decoder architecture (Vaswani et al., 2017) and scales the training data to 680k hours of proprietary labelled audio. [^1] Their latest large-v3 version uses 1M hours of labelled audio and 4M hours of pseudo-labelled audio. Despite its strong performance on ASR, ST, and LID, the full development pipeline including training data details and training codebase is not publicly available. A recent project OWSM aims to reproduce Whisper-style training using public data and open-source toolkits to promote transparency and open science in this field (Peng et al., 2023e). The latest OWSM v3.1 models (Peng et al., 2024) employ E-Branchformer (Kim et al., 2023) as the encoder and Transformer as the decoder, which is trained with a joint ASR CTC loss (Kim et al., 2017). Although OWSM has promising results using public corpora, it still follows the encoder-decoder architecture, which can be slow and unstable at inference time.
@@
CTC or Transducer. Another line of research proposes to utilize CTC (Graves et al., 2006) or Transducer (Graves, 2012) for ASR. Google’s USM (Zhang et al., 2023) provides generic ASR models, which are first pre-trained on 12M hours of unlabelled audio and then fine-tuned on proprietary labelled data with CTC or Transducer. Meta’s MMS (Pratap et al., 2023) pre-trains a wav2vec 2.0 model (Baevski et al., 2020) on massively multilingual data and then fine-tunes it with CTC on labelled ASR data covering over 1k languages. These models employ CTC only for ASR. In our OWSM-CTC, we propose a single CTC-based encoder-only model for ASR, ST and LID. Our supported tasks are more similar to Whisper-style models.
@@ 2.2 Efficient speech models
Efficient architectures. Better network architectures can also improve efficiency, including attention with linear complexity (Beltagy et al., 2020; Wang et al., 2020b; Tay et al., 2023) and sequence length reduction (Burchi and Vielzeuf, 2021; Kim et al., 2022; Nawrot et al., 2023; Rekesh et al., 2023). In this work, we do not modify the attention, but we use larger downsampling in CNN to reduce the sequence length. More details are in Appendix A.2 and B.1.
@@ 2.3 CTC-based speech models
Although CTC assumes a monotonic alignment between input and output, it is promising for ST due to the reordering capability of self-attention (Inaguma et al., 2021; Chuang et al., 2021).
Conventional CTC models are typically designed for a specific task or language. It remains under-explored whether such approaches can be scaled to multilingual and multi-task scenarios. This work proposes a novel encoder-only speech foundation model based on multi-task self-conditioned CTC. This single model performs well in multilingual ASR, ST and LID.
@@ 3.1 Overall architecture
Figure 2 shows the architecture of OWSM-CTC. Its main component is a speech encoder, which takes speech features as input and predicts the spoken language as well as the ASR or ST hypothesis using CTC. To mimic Whisper-style models, which condition text generation on an optional text prompt (Radford et al., 2023; Peng et al., 2023e, 2024), we employ a separate Transformer encoder to process the prompt and inject the output to the main model through cross-attention. Then, the model can potentially attend to the text prompt when generating text.
@@ 3.3 Prompt encoder
Our training data is a mixture of public ASR and ST datasets. Some of them provide unsegmented long audio, but the others only release segmented short audio. At training time, if the sample does not have a previous sentence, we will use <na>. Otherwise, we use either <na> or the previous sentence as the prompt according to 0.5 probability. Section 4.6 shows that OWSM-CTC can leverage the prompt’s information when necessary.
@@
Table 1: Summary of model size, training data and training cost measured on NVIDIA A100 GPU (40GB).
@@ 4.1 Experimental setups
Evaluation. We fairly compare our encoder-only OWSM-CTC with the previously released encoder-decoder OWSM v3.1 models (Peng et al., 2024) since they are trained on the same data. We also show the results of Whisper under the same decoding setup for reference, but we note that they are not comparable with ours due to completely different training data. By default, short-form audio without any text prompt is used, but we also evaluate the long-form ASR performance in Section 4.5 and investigate the effect of text prompt in Section 4.6.
@@
Table 2: Spoken language identification results on the FLEURS test set. Bold: the best result. Underlined: our OWSM-CTC outperforms OWSM v3.1 medium.
@@ 4.2 Language identification
Table 2 presents the LID results on the FLEURS test set. Our OWSM-CTC achieves a top-1 accuracy of 87.6%, outperforming the other encoder-decoder models by a large margin. This is likely because spoken LID requires a powerful encoder to extract useful information from the input audio. Our encoder-only model is especially suitable for this type of task.
@@
Table 3: WER % ($\downarrow$) of English ASR. Speed-up ($\uparrow$) is measured using the average decoding time. Whisper is trained on 438k hours of English audio, whereas OWSM v3.1 and our OWSM-CTC are trained on only 73k hours. Bold: the best result. Underlined: our OWSM-CTC outperforms OWSM v3.1 medium.
@@
Table 4: Multilingual ASR results. CER % ($\downarrow$) is shown for Chinese (zh), Korean (ko) and Japanese (ja), while WER % ($\downarrow$) is shown for the others. Data sizes are in thousand hours. Bold: the best result. Underlined: our OWSM-CTC outperforms OWSM v3.1 medium.
@@ 4.3 Speech recognition
Table 3 presents word error rates (WERs) on multiple English ASR test sets. Following Peng et al. (2023e, 2024), we leverage greedy decoding and apply the Whisper English text normalizer before scoring. We record the average decoding time across all English test sets on NVIDIA A40 GPU and calculate the relative speed-up. Results show that our non-autoregressive OWSM-CTC generally has comparable WERs with the autoregressive OWSM v3.1 medium (average: 7.9 vs. 7.7), both of which have 1B parameters. However, OWSM-CTC achieves 3.63x speed-up due to parallel decoding. Notably, OWSM-CTC is even faster than OWSM v3.1 base, which has only 100M parameters, and our WERs are much lower (average: 7.9 vs. 12.1). Compared to Whisper models trained on significantly more data, our OWSM-CTC is still competitive in many cases, and our inference is much faster. These results demonstrate that OWSM-CTC achieves an excellent trade-off between recognition accuracy and inference efficiency.
Table 4 shows the results of multilingual ASR. We perform greedy decoding and apply the Whisper basic text normalizer before scoring. Our OWSM-CTC is slightly worse than OWSM v3.1 in terms of the average WER/CER (16.2 vs. 15.2). For European languages in MLS (Pratap et al., 2020), OWSM-CTC generally falls behind. But for East Asian languages like Chinese, Japanese and Korean, OWSM-CTC is on par with or better than OWSM v3.1 medium. This difference might be related to the training data size and tokenization.
Src Lang. de es fr ca Average ($\uparrow$) Speed-up ($\uparrow$)
@@
base 11.4 19.2 13.1 9.7 13.4 1.84x
small 25.0 32.8 26.4 21.7 26.5 1.54x
medium 33.6 39.7 34.4 29.2 34.2 0.84x
@@
medium 21.1 28.2 27.7 23.7 25.2 3.35x
Table 5: BLEU ($\uparrow$) of X-to-En ST on CoVoST-2. Speed-up is measured using average decoding time. Data sizes are in thousand hours. Bold: the best result. Underlined: our OWSM-CTC outperforms OWSM v3.1 medium.
data size 14.0 0.4 13.7 0.8 0.4 0.4 0.9 0.9 0.4 0.4 0.4 0.4 1.0 0.4 0.4
@@
Table 6: BLEU ($\uparrow$) of En-to-X ST on CoVoST-2. Speed-up is measured using the average decoding time across all 15 directions. Data sizes are in thousand hours. Bold: the best result. Underlined: our OWSM-CTC outperforms OWSM v3.1 medium. Note that Whisper (Radford et al., 2023) does not support En-to-X translation.

### v2

@@ Abstract
There has been an increasing interest in large speech models that can perform multiple tasks in a single model. Such models usually adopt an encoder-decoder or decoder-only architecture due to their popularity and good performance in many domains. However, autoregressive models can be slower during inference compared to non-autoregressive models and also have potential risks of hallucination. Though prior studies observed promising results of non-autoregressive models for certain tasks at small scales, it remains unclear if they can be scaled to speech-to-text generation in diverse languages and tasks. Inspired by the Open Whisper-style Speech Model (OWSM) project, we propose OWSM-CTC, a novel encoder-only speech foundation model based on Connectionist Temporal Classification (CTC). It is trained on 180k hours of public audio data for multilingual automatic speech recognition (ASR), speech translation (ST), and language identification (LID). Compared to encoder-decoder OWSM, our OWSM-CTC achieves competitive results on ASR and up to 24% relative improvement on ST, while it is more robust and 3 to 4 times faster for inference. OWSM-CTC also improves the long-form ASR result with 20x speed-up. We will publicly release our code, pre-trained model, and training logs to promote open science in speech foundation models. [^1] https://github.com/espnet/espnet
@@ 1 Introduction
111122223333444488881010101012121212mediumbaseSpeed-up ($\rightarrow$)WER ($\leftarrow$)OWSM v3.1 seriesOWSM-CTC (ours)
@@
1111222233334444101010101515151520202020252525253030303035353535mediumbaseSpeed-up ($\rightarrow$)BLEU ($\rightarrow$)OWSM v3.1 seriesOWSM-CTC (ours)
@@
1111222233334444555510101010151515152020202025252525mediumbaseSpeed-up ($\rightarrow$)BLEU ($\rightarrow$)OWSM v3.1 seriesOWSM-CTC (ours)
@@
The great success of large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023; Anil et al., 2023b) has sparked a growing interest in developing foundation models in various modalities. Recent studies have explored different approaches towards multilingual and multi-tasking speech foundation models (Radford et al., 2023; Zhang et al., 2023; Pratap et al., 2023; Rubenstein et al., 2023; Barrault et al., 2023; Peng et al., 2023e). OpenAI Whisper (Radford et al., 2023) is a series of Transformer encoder-decoder models trained on 680k hours of proprietary labeled audio. Whisper achieves strong results in multilingual automatic speech recognition (ASR), any-to-English speech translation (ST), and spoken language identification (LID). Although it shows the effectiveness of large-scale (weakly) supervised pre-training, the full development pipeline, including training data details, is not publicly accessible. Recent works have developed Open Whisper-style Speech Models (OWSM) (Peng et al., 2023e, 2024) with the aim of reproducing Whisper-style training using public data and open-source toolkits. However, Whisper and OWSM adopt the encoder-decoder architecture, which generates text tokens given speech in an autoregressive manner. They might hallucinate during inference, and the speed can be slow. Other models with decoder-only architectures, like AudioPaLM (Rubenstein et al., 2023) and VioLA (Wang et al., 2023b), could suffer from the same issues due to autoregressive decoding.
@@
Another type of work like Google USM (Zhang et al., 2023) and Meta MMS (Pratap et al., 2023) uses non-autoregressive models with Connectionist Temporal Classification (CTC) Graves et al. (2006), but these CTC-based models are designed for ASR only. Prior studies have also achieved promising results of CTC models for ST only, but they mainly focus on specific language pairs at much smaller scales (Inaguma et al., 2021; Chuang et al., 2021; Xu et al., 2023). Some of them employ additional decoders (Inaguma et al., 2021; Yan et al., 2023) or cross-attention layers (Xu et al., 2023), making the model more complicated.
@@
A natural question now arises: Can we build a non-autoregressive encoder-only model for speech-to-text generation in diverse languages and multiple tasks like Whisper/OWSM? This research problem has become increasingly important in the era of LLMs because large-scale pre-trained speech encoders can serve as an adapter between the speech and text modalities (Gong et al., 2023; Wang et al., 2023a), providing a promising avenue towards general-purpose multi-modal foundation models (Anil et al., 2023a).
@@
In this work, we propose OWSM-CTC, a novel encoder-only speech foundation model based on multi-task self-conditioned CTC to imitate OWSM’s multilingual ASR, any-to-any ST, and LID functionalities. Following previous encoder-decoder OWSM v3.1 models (Peng et al., 2024), we train a 1B OWSM-CTC model using 180k hours of public data covering 151 languages. Extensive evaluations show that our OWSM-CTC exhibits strong performance and efficiency. Compared to the 1B OWSM v3.1 medium model, OWSM-CTC achieves comparable performance for ASR and superior performance for various ST directions (up to 24% relative improvement) while being more robust and showing 3 to 4 times inference speed-up. OWSM-CTC also improves the WER for long-form ASR and can be 20 times faster due to batched parallel decoding. OWSM-CTC further outperforms the other baseline models on LID. Our code, pre-trained model weights, and training logs will be publicly released to facilitate the development of large speech models.
@@ 2.1 Speech foundation models
Attention-based encoder-decoder. OpenAI Whisper (Radford et al., 2023) adopts the standard Transformer encoder-decoder architecture (Vaswani et al., 2017) and scales the training data to 680k hours of proprietary labeled audio. [^2] Their latest large-v3 version uses 1M hours of labeled audio and 4M hours of pseudo-labeled audio. However, the complete pipeline for model development, including training data details and training code, is not publicly available. A recent project, OWSM, aims to reproduce Whisper-style training using public data and open-source toolkits to promote transparency and open science in this field (Peng et al., 2023e). The latest OWSM v3.1 models (Peng et al., 2024) employ E-Branchformer (Kim et al., 2023) as the encoder and Transformer as the decoder, which are trained with a joint ASR CTC loss (Kim et al., 2017). Although OWSM has promising results using public corpora, it still follows the encoder-decoder architecture, which can be slow and unstable at inference time.
@@
CTC or Transducer. Another line of research proposes to utilize CTC (Graves et al., 2006) or Transducer (Graves, 2012) for ASR. Google USM (Zhang et al., 2023) provides generic ASR models that are first pre-trained on 12M hours of unlabeled audio and then fine-tuned on proprietary labeled data with CTC or Transducer. Meta MMS (Pratap et al., 2023) pre-trains a wav2vec 2.0 model (Baevski et al., 2020) on massively multilingual data and then fine-tunes it with CTC on labeled ASR data covering over 1k languages. These models employ CTC only for ASR. In our OWSM-CTC, we propose a single CTC-based encoder-only model for ASR, ST, and LID. Our supported tasks are more similar to Whisper-style models.
@@ 2.2 Efficient speech models
Efficient architectures. Better network architectures can also improve efficiency, including attention with linear complexity (Beltagy et al., 2020; Wang et al., 2020b; Tay et al., 2023) and sequence length reduction (Burchi and Vielzeuf, 2021; Kim et al., 2022; Nawrot et al., 2023; Rekesh et al., 2023). In this work, we do not modify the attention but use larger downsampling in the convolution module to reduce the sequence length. More details are in Appendix A.2 and B.1.
@@ 2.3 CTC-based speech models
Although CTC assumes a monotonic alignment between input and output, it can be used for ST with the reordering capability of self-attention (Inaguma et al., 2021; Chuang et al., 2021).
Conventional CTC models are typically designed for a specific task or language. It remains under-explored whether such approaches can be scaled to multilingual and multi-task scenarios. This work proposes a novel encoder-only speech foundation model based on multi-task self-conditioned CTC. This single model performs well in multilingual ASR, ST, and LID.
@@ 3.1 Overall architecture
Figure 2 shows the architecture of OWSM-CTC. Its main component is a speech encoder, which takes speech features as input and predicts the spoken language as well as the ASR or ST hypothesis using CTC. To mimic Whisper-style models that condition text generation on an optional text prompt (Radford et al., 2023; Peng et al., 2023e, 2024), we employ a separate Transformer encoder to process the prompt and inject the output to the main model through cross-attention. Then, the model can potentially attend to the text prompt when generating text.
@@ 3.3 Prompt encoder
Our training data is a mixture of public ASR and ST datasets. Some of them provide unsegmented long audio, but the others only release segmented short audio. At training time, if a sample does not have a previous sentence, we will use <na>. Otherwise, we use either <na> or the previous sentence as the prompt according to 0.5 probability. Section 4.6 shows that OWSM-CTC can leverage the prompt’s information when necessary.
@@
Table 1: Summary of model size, training data, and training cost measured on an NVIDIA A100 GPU (40GB).
@@ 4.1 Experimental setups
Evaluation. We fairly compare our encoder-only OWSM-CTC with the previously released encoder-decoder OWSM v3.1 models (Peng et al., 2024) since they are trained on the same data. We also show the results of Whisper under the same decoding setup for reference, but we note that they are not comparable with ours due to completely different training data. By default, short-form audio without any text prompt is used, but we also evaluate the long-form ASR performance in Section 4.5 and investigate the effect of text prompts in Section 4.6.
@@
Table 2: Spoken LID results on the FLEURS test set.
@@ 4.2 Language identification
Table 2 presents the LID results on the FLEURS test set (Conneau et al., 2023). Our OWSM-CTC achieves a top-1 accuracy of 87.6%, outperforming the other encoder-decoder models by a large margin. This is likely because spoken LID requires a powerful encoder to extract useful information from the input audio. Our encoder-only model is especially suitable for this type of task.
@@
Table 3: WER % ($\downarrow$) of English ASR. Speed-up ($\uparrow$) is based on average decoding time. Whisper is trained on 438k hours of English audio, whereas OWSM v3.1 and our OWSM-CTC are trained on only 73k hours. Results of Whisper large-v2 and OWSM v3.1 medium with beam search are shown in gray, which are not comparable with the others due to different model sizes or decoding configurations. Bold: the best result. Underlined: OWSM-CTC outperforms OWSM v3.1 medium.
@@
Table 4: Multilingual ASR results. CER % ($\downarrow$) is shown for Chinese (zh), Korean (ko) and Japanese (ja), while WER % ($\downarrow$) is shown for the others. Data sizes are in thousand hours. Results of Whisper large-v2 and OWSM v3.1 medium with beam search are shown in gray, which are not comparable with the others. Bold: the best result. Underlined: OWSM-CTC outperforms OWSM v3.1 medium.
@@ 4.3 Speech recognition
Table 3 presents word error rates (WERs) on nine English ASR test sets. Following Peng et al. (2023e, 2024), we leverage greedy decoding and apply the Whisper English text normalizer before scoring. [^3] We also report the results of Whisper large-v2 and OWSM v3.1 medium with beam search in gray for reference, but they are not comparable with the others due to different model sizes or decoding configurations. This applies to other tables as well. We record the average decoding time across all English test sets on an NVIDIA A40 GPU and calculate the relative speed-up. Results show that our non-autoregressive OWSM-CTC generally has comparable WERs with the autoregressive OWSM v3.1 medium (average: 7.9 vs. 7.7), both of which have 1B parameters. However, OWSM-CTC achieves 3.63x speed-up due to parallel decoding. Notably, OWSM-CTC is even faster than OWSM v3.1 base, which has only 100M parameters, and our WERs are much lower (average: 7.9 vs. 12.1). Compared to Whisper models trained on significantly more data, our OWSM-CTC is still competitive in many cases, and our inference is much faster. These results demonstrate that OWSM-CTC achieves an excellent trade-off between recognition accuracy and inference efficiency.
Table 4 shows the results of multilingual ASR. We perform greedy decoding and apply the Whisper basic text normalizer before scoring. Our OWSM-CTC is slightly worse than OWSM v3.1 in terms of the average WER/CER (16.2 vs. 15.2). For European languages in MLS (Pratap et al., 2020), OWSM-CTC generally falls behind. But for East Asian languages like Chinese, Japanese, and Korean, OWSM-CTC is on par with or better than OWSM v3.1 medium. This difference might be related to the training data size and tokenization.
Src Lang. de es fr ca Ave. ($\uparrow$) Speed-up ($\uparrow$)
@@
base 11.0 18.9 13.2 9.9 13.3 1.84x
small 23.9 31.8 26.1 21.4 25.8 1.54x
medium 32.0 37.3 33.4 28.8 32.9 0.84x
large-v2 35.2 39.7 35.7 31.2 35.5 0.48x
@@
medium 20.7 27.9 27.5 24.2 25.1 3.35x
Table 5: BLEU ($\uparrow$) of X-to-En ST on CoVoST-2. Data sizes are in thousand hours. Results of Whisper large-v2 and OWSM v3.1 medium with beam search are shown in gray, which are not comparable with the others. Bold: the best result. Underlined: OWSM-CTC outperforms OWSM v3.1 medium.
Tgt Lang. de ca zh fa et mn tr ar sv lv sl ta ja id cy Ave. ($\uparrow$) Speed-up ($\uparrow$)
data size 14.0 0.4 13.7 0.8 0.4 0.4 0.9 0.9 0.4 0.4 0.4 0.4 1.0 0.4 0.4 - -
@@
medium 26.7 24.0 32.9 9.9 11.4 6.2 7.9 8.3 24.5 10.0 14.2 0.1 20.4 22.6 20.6 16.0 4.20x
p-value 0.006 0.001 0.001 0.001 0.001 0.001 0.145 0.001 0.001 0.001 0.001 0.031 0.001 0.001 0.001 - -
Table 6: BLEU ($\uparrow$) of En-to-X ST on CoVoST-2. Data sizes are in thousand hours. Note that Whisper does not support En-to-X translation. The p-values are computed by comparing OWSM-CTC against OWSM v3.1 medium using the Paired Significance Test in SacreBLEU (Post, 2018). Results of OWSM v3.1 medium with beam search are shown in gray, which are not comparable with the others.